form ammsvúa as. Calaul ate tle volune on annonia 945
Answer #1

reaction is

N2 + 3 H2 ................> 2 NH3

one mole N2 react with 3 mole H2 to form 2 mole NH3.

mole of N2 = mass / molar mass = 28 g / 28 g / mole = 1.0 mole.

N2 is the limiting reactant.

thus

mole of NH3 produce = 2 * 1.0 = 2.0 mole.

ideal gas equation is

PV = nRT

1 atm * V = 2.0 mole * 0.0821 L-atm / mole-K * 303 K

or

V = 49.75 L

volume of NH3 gas = 49.75 L